. Prove that if n is a perfect square, then n + 2 is not a



perfect square.




1
Expert's answer
2022-04-18T01:52:00-0400

Let "n=k^2," where "k" is some integer number.

Let's find the nearest to the "n" number that is a perfect square and is greater than "n".

Obviously, this number is "(k+1)^2=k^2+2k+1=n+2k+1" and it is bigger than "n+2" when "k\ge1". So, "\sqrt{n+2}" should lie between "k" and "k+1" and it is not an integer number.